Question 2 Cups of coffee and donuts are complements. Both have inelastic demand. A hurricane destroys half the coffee bean crop. Use appropriately labelled diagrams to answer the following questions. a. What happens to the price of coffee beans? b. What happens to the price of a cup of coffee? What happens to total expenditure on cups of coffee? c. What happens to the price of donuts? What happens to total expenditure on donuts? Question 3 Consider public policy aimed at smoking. a. Studies indicate that the price elasticity of demand for cigarettes is about 0.4. If a pack of cigarettes currently costs $2 and the government wants to reduce smoking by 20 percent, by how much should it increase the price? b. If the government permanently increases the price of cigarettes, will the policy have a larger effect on smoking one year from now or five years from now? INTRODUCTION TO ECONOMICS I ECON 101 2 c. Studies also find that teenagers have a higher price elasticity of demand than do adults. Why might this be true?


1
Expert's answer
2021-03-14T19:05:30-0400

2(a)The price of of coffee beans increases due to the reduced supply.When a hurricane destroys half of the crop, the supply of coffee beans decreases from S1 to S2, the price of coffee beans increases, and the quantity decreases.



2(b). The price of a cup of coffee will increase due to the decreased supply of coffee.When the price of coffee beans increases, the quantity supply of cups of coffee decreases, the price of a cup of coffee increases, and the quantity decreases.



The total expenditure of coffee will rise. This is because for goods with inelastic demand ,increase in price of a good raises the total expenditure on that good.

2(c)The price of donuts decreases.When the price of coffee increases and the quantity demanded of coffee decreases, consumers demand fewer donuts because coffee and donuts are complements. The demand for donuts will decrease from D2 to D2 as shown below , lowering the donuts price.

The total expenditure of donuts will decrease. Donuts is an inelastic good and decrease in price will result to decreased expenditure on donuts.

3(a)price elasticity of demand"=\Delta Q/\Delta P"

"=0.4=20\/p"

"\Delta P=50" %

"2-x\/2=-50\/100"

"2-x\/2=-0.5"

"2-x=-0.5*2"

"x=2+1"

"x=3"

The government should increase the price to $3 from $2

3(b)The policy will have a larger effect on smoking five years from now than in the short run where demand is elastic. This is because goods tend to have more elastic demand over long time. In the short run the consumers might not respond to the price change but in the long run the consumers will reduce their consumption due to high prices.

3(c) The teenagers are responsive to price changes because of they have limited financial resources to purchase cigarettes compared to adults. Teenagers are new to smoking compared to the adults who are addicts and find it difficult to reduce smoking even when the price increases.